OpenStudy (anonymous):

Jodi went 54 miles on 3.5 gallons of gas, and she wants to know how far she can go on 9 gallons of gas. Which proportion would be useful to solve the problem?

OpenStudy (imstuck):

Do the proportion like this:\[\frac{ m }{ g }=\frac{ m }{g }\]where m is the number of miles sh can go on g, the gallons of gas. You have that Jodi can go 54 miles on 3.5 gallons, so that's your first proportion:

OpenStudy (imstuck):

\[\frac{ 54 }{ 3.5 }=\frac{ m }{ g }\]

OpenStudy (imstuck):

You want to know how far she can go on 9 gallons of gas, so the 9 goes into the second proportion in the g position:

OpenStudy (imstuck):

\[\frac{ 54 }{3.5 }=\frac{ m }{ 9 }\]To solve this, you cross multiply: 54(9)=3.5m

OpenStudy (imstuck):

486=3.5m

OpenStudy (imstuck):

Can you solve that? When you do, the value you get for m is approximately how far she can go on 9 gallons of gas.